Max grabs three marbles from a bag containing five white marbles and six black marbles. What...

  1. Max grabs three marbles from a bag containing five white marbles and six black marbles. What is the probability that exactly two of the marbles are black?
Answer #1

Please note

Total Marbles in the bag = 5 white + 6 black = 11

Probability = Favorable Outcomes / Total Outcomes

Favorable outcomes = Choosing 2 black marbles out of 6 and 1 white out of 5 = 6C2 + 5C1 = (6 * 5 / 2) + 5 = 20

Total outcomes = Choosing 3 marbles out of 11 = 11C3 = 11 * 10 * 9 / (3 * 2 * 1) = 165

Therefore the required probability = 20 / 165 = 4 / 33 = 0.1212